OpenStudy (anonymous):

BC trucking company charges $56(P) to deliver 8 cartons(c) to a nearby town, and $78 to deliver 24 cartons Find a linear equation for these data points using (c,f then use this equation to find how much BC company charges to deliver 72 cartons

OpenStudy (anonymous):

let y be the company charge and x be the number of cartons to be delivered. now we have 2 points, (8,56) and (24,78). first, find the slope, m=(y2-y1)/(x2-x1)=32/16=2 substitute into general equation of a straight line, y=2x+c which c is the y-intercept. to find the value of c, pick a point and plug in the value into the equation. i.e. we pick (8,56) which x=8 and y=56. y=2x+c, 56=2(8)+c, 40=c. now you have a straight line y=2x+40. to know how much the charger of delivering 72 cartons, just plug in x=72 into the equation of straight line :)